OpenStudy (anonymous):

The difference of the ages of two very young children is 2. Three times the younger child plus two times the older child is 9. Find the age of each child. (Set up equations.)

OpenStudy (anonymous):

1 and 3

OpenStudy (anonymous):

can you help me set up an equation for it

OpenStudy (anonymous):

um i guess

OpenStudy (anonymous):

N will be younger child N+2 will be older child

OpenStudy (anonymous):

would the first equation be something like x-y=2 or something like that?

OpenStudy (anonymous):

3N + 2(N+2) = 9

OpenStudy (anonymous):

do you understand that ?

OpenStudy (anonymous):

why are you multiplying the 2?

OpenStudy (anonymous):

two times the older child

OpenStudy (anonymous):

then why are you adding 2 in parentheses?

OpenStudy (anonymous):

N will be younger child... this is the young boy N+2 will be older child this is the older boy... *Now I added 2 in parentheses because the older boy is the same age as the younger but 2 years older so... you add 2

OpenStudy (anonymous):

ok what about the first part of the problem;The difference of the ages of two very young children is 2

OpenStudy (anonymous):

you need an equation for that right also?

OpenStudy (anonymous):

wouldnt it be something like x-y=2

OpenStudy (anonymous):

??

OpenStudy (anonymous):

are you still here?